OpenStudy (anonymous):

Let u = <-5, 1>, v = <7, -4>. Find 9u - 6v. A. <18, 18> B. <-87, 33> C. <-36, -18> D. <-3, -15>

OpenStudy (anonymous):

are you in connections?

OpenStudy (anonymous):

ikr lol i am

OpenStudy (anonymous):

same, and i'm working on this same exact thing...

OpenStudy (anonymous):

not me lol im in 7th gradde...

OpenStudy (anonymous):

what grade is this

OpenStudy (anonymous):

this is algebra 1. but i'm a sophomore.

OpenStudy (anonymous):

Cool i have to do Surface Area and Volume right now i already did algreba

OpenStudy (anonymous):

Live lesson today in 30 minutes .

OpenStudy (anonymous):

how did you get that

OpenStudy (anonymous):

is there any special meaning of the <> notation ..?

OpenStudy (anonymous):

9u= 9*-5=-45; 6v=6*7=42 9u-6v=-87 similarly calculate for the 2nd terms in the <>

OpenStudy (anonymous):

thanks